OpenStudy (anonymous):

The equation 2x2 - 2x - 12 = 0 is factored. Each factor is set equal to zero. What are these two equations? A. 2x - 3 = 0, x + 4 = 0 B. 2x - 4 = 0, x + 3 = 0 C. 2x + 6 = 0, x - 2 = 0 D. 2x - 6 = 0, x + 2 = 0

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Ok so first we factor it yes? I got \[(x-3)(2x+4)\] With this we can set BOTH factors of this equation to 0, cause remember ANYTHING times 0 is always 0 so if one of these factors is zero the WHOLE THING is 0. \[x-3 = 0\] \[x=3\] ---- \[2x+4 = 0\] \[x=4/2\] \[x=2\]

OpenStudy (anonymous):

That's my only problem is that's not an option for what you gave... so hmm

OpenStudy (anonymous):

Double check B see if you accidentally flipped the signs when typing it out?

OpenStudy (anonymous):

nope not accident

OpenStudy (anonymous):

Well this is funny you can factor it both ways so... D is correct because \[2x^2 - 2x - 12 = (2x-6)(x+2)\] So there are zeroes at 3 and 2 just like above. That's funny never seen that before.